OpenStudy (anonymous):

Write in the form a + bi. -3/(2 + i)

OpenStudy (saifoo.khan):

-1.2 + 0.6i

OpenStudy (anonymous):

How did you do that?

OpenStudy (saifoo.khan):

Calc.

OpenStudy (anonymous):

$$ \dfrac{-3}{2+i}=\dfrac{-3(2-i)}{(2-i)(2+i)}=\dfrac{-6+3i}{2^2-i^2}=\dfrac{-6+3i}{4-(-1)}=\dfrac{-6+3i}{5}=\dfrac{-6}{5}+\dfrac{3}{5}i=-1.2+0.6i$$

OpenStudy (anonymous):

Gone off the edge of the page but right idea.

OpenStudy (anonymous):

Thank you!!! Why do we times by the negative version of the denominator?

OpenStudy (anonymous):

It's called the conjugate. Similar to when u simplify radicals, the idea is that u want to get rid of i's on the bottom. As u can see, the things cancel out nicely and leave u with a number on the bottom when u do this.